Mr. wasif spent 1/3 of his income on his housing and 2/5 of his income on his daily expenditure . He saved the rest. Last Year Mr. Wasif saved Rs 24,600 . What was gthe toatl income of Mr wasif last year??
Answer by Theo(13342)

he spent 1/3 of his income on housing and 2/5 of his income on his daily expenditures.
he saved the rest, which was equal to 24600.
what was his total income?

total that he spent was 1/3 + 2/5 of his income.
1/3 * 5/5 = 5/15
2/5 * 3/3 = 6/15
total that he spent was 5/15 + 6/15 = 11/15 of his income.
that means he saved 4/15 of his income.

let x = his income.
4/15 * x = 24,600
multiply both sides of this equation by 15/4 to get:
4/15 * 15/4 * x = 24,600 * 15/4
solve for x to get:
x = 92,250.

to check do the following/

he spent 1/3 of his income on housing.
that becomes 1/3 * 92,250 = 30,750

he spent 2/5 of his income on his daily expenditure.
that becomes 2/5 * 92,250 = 36,900

add those up to get total expenditures of 67,750

subtract that from 92,250 to get 24,600

the solution looks good.

his total income last year was 92,250.